How would I factor 15x^2-4x-4=0?
Answer by Alan3354(69443)

How would I factor 15x^2-4x-4=0?
==================
You find a pair of factors of 15, and a pair of 4 that give the product.
It's trial and error, having someone do it for you doesn't accomplish anything.
-----
(ax + b)*(cx + d) = 15x^2-4x-4
a*c = 15
b*d = -4